Antioch Extraction, which mines ore in Montana, uses a calendar year for both financial-reporting and tax purposes. The following selected costs were incurred in December, the low point of activity, when 1,450 tons of ore were extracted:

Straight-line depreciation $ 33,500
Charitable contributions* 9,500
Mining labor/fringe benefits 261,000
Royalties 147,250
Trucking and hauling 255,175

*Incurred only in December.

Peak activity of 2,750 tons occurred in June, resulting in mining labor/fringe benefit costs of $495,000, royalties of $244,750, and trucking and hauling outlays of $330,175. The trucking and hauling outlays exhibit the following behavior:

Less than 1,450 tons $ 217,675
From 1,450–1,949 tons 255,175
From 1,950–2,449 tons 292,675
From 2,450–2,949 tons 330,175

Antioch uses the high-low method to analyze costs.

Required:

1. Classify the five costs listed in terms of their behavior: variable, step-variable, committed fixed, discretionary fixed, step-fixed, or semivariable.

2. Calculate the total cost for next February when 1,750 tons are expected to be extracted.

3-a. Is hauling 1,450 tons with respect to Antioch’s trucking/hauling cost behavior cost-effective?

3-b. If the company plans to extract 1,450 tons, at what number of tons can cost-effectiveness be achieved?

4. Distinguish between committed and discretionary fixed costs. If Antioch were to experience severe economic difficulties, which of the two types of fixed costs should management try to cut?

5. Speculate as to why the company’s charitable contribution cost arises only in December.
Answer #1

Part 1)

The classification is provided as below:

Cost Classification
Straight-line depreciation Committed Fixed Cost
Charitable Contributions Discretionary Fixed Cost
Mining Labor/Fringe Benefits Variable
Royalties Semi-Variable Cost
Trucking and Hauling Step-Fixed Cost

______

Notes:

1) The per-ton mining labor/fringe benefits cost at two different levels (as shown below) remains same, indicating that it is a variable cost that increases with increase in activity level/production volume.

Per-Ton Mining Labor/Fringe Benefits Cost at 1,450 Tons = 261,000/1,450 = $180 per ton

Per-Ton Mining Labor/Fringe Benefits Cost at 2,750 Tons = 495,000/2,750 = $180 per ton

2) Royalty comprises of both fixed and variable cost (as shown below). Therefore, its variable part will increase with an increase in activity level/production volume while the fixed part will remain constant irrespective of increase/decrease in activity level/production volume.

Variable Portion of Royalty Cost = (Total Royalty Cost at 2,750 Tons - Total Royalty Cost at 1,450 Tons)/(2,750 - 1,450) = (244,750 - 147,250)/(2,750 - 1,450) = $75 per ton

Fixed Portion of Royalty Cost:

December (1,450 Tons) June (2,750 Tons)
Total Royalty Cost 147,250 244,750
Less Variable Portion of Royalty Cost 108,750 (1,450*75) 206,250 (2,750*75)
Fixed Portion of Royalty Cost $38,500 $38,500

______

Part 2)

The total cost for next February when 1,750 tons are expected to be extracted is arrived as below:

Straight Line Depreciation 33,500
Charitable Contributions 0
Mining Labor/Fringe Benefits (1,750*180) 315,000
Royalties:
Fixed 38,500
Variable (1,750*75) 131,250
Trucking and hauling (Refer to table) 255,175
Total Cost $773,425

Answer for Part 2) is $773,425.

______

Part 3-a)

No, hauling 1,450 tons with respect to Antioch’s trucking/hauling cost behavior is not effective. It is because the company would incur $255,175 between the activity level/production volume of 1,450 -1,949 tons while it would spend $217,675 at an activity level/production volume of less than 1,450 units. Therefore, the company would be better off at an activity level/production volume of 1,449 tons as it would save $37,500 (255,175 - 217,675) in trucking/hauling cost.

______

Part 3-b)

As mentioned above (in Part 3-a), the company would be able to achieve cost-effectiveness at 1,449 tons if it plans to extract 1,450 tons. At an activity level/production volume of 1,449 units, the company would incur $217,675. However, at an activity level/production volume of 1,450 tons, the company would spend $255,175. In other words, in case of step-fixed cost, the company is able to achieve cost effectiveness it it operates at the right-most level of the step (exactly before the jump in cost).

Answer for Part 3-b) is 1,449 tons

______

Part 4)

Committed fixed costs arise as a result of use of company's resources/facilities such as depreciation resulting from use of plant and equipment or rent paid for office building or salaries paid to managers/employees for their services and so on. Discretionary fixed costs, on the other hand, are incurred for some specific purpose such as charity, research and development, advertising and so on. The company is not under any obligation to spend money on such activities. It may, at its own discretion, incur such expenses as and when required.

Since, the company is under no obligation to spend money on discretionary fixed costs, it should try to cut discretionary fixed costs if it experiences severe economic difficulties. It is because the company can incur such costs as and when there is an absolute need or company has surplus funds. Further, elimination of such costs may not have any adverse impact on the business operations in the long run. For instance, the company may decide not to make charitable contributions in a year if its profits declined because of economic crisis.

______

Part 5)

Company's charitable contribution cost arises only in December because of surplus funds that may be available at the year end (Antioch Extraction uses calendar year for both financial-reporting and tax purposes). Company's management may decide to donate some amount of money at the closure of business year as a part of its social objectives or to reduce its tax liability. Contribution to charities are tax-deductible which help in reducing taxable income and consequently the tax liability of the company.